OpenStudy (anonymous):

y + 3x = 0 y − 6 = −3x^2 how do i solve

OpenStudy (anonymous):

you can write the first equation as \(y=-3x\) and then solve \[-3x-6=-3x^2\]

OpenStudy (anonymous):

add \(3x^2\) to ge t \[3x^2-3x-6=0\] and then factor as \[3(x^2-x-2)=0\]

OpenStudy (anonymous):

that will help me get this ?

OpenStudy (anonymous):

yes it will

OpenStudy (anonymous):

solve \[3(x-2)(x+1)=0\] and so there are two solution for \(x\) namely \(2\) or \(-1\) now we have to find \(y\)

OpenStudy (anonymous):

which is not hard, since we already know \(y=-3x\)

OpenStudy (anonymous):

so if \(x=2\) then \(y=-3\times 2=-6\) and one solution is \((2,-6)\)

OpenStudy (anonymous):

if \(x=-1\) then \(y=-3\times (-1)=3\) so another solution is \((-1,3)\)

OpenStudy (anonymous):

so much work! thanks babe!